OpenStudy (anonymous):

if you were to use the substitution method to solve the following system, choose the new system of equations that would result if x was isolated in the first equation. x-6y+3z=0 2x-5+8z=7 x+7y-4z=13 a) -17y+14z=7 y-z=13 b) 7y+2z=7 13y-7z=13 c) 13y+7z=-13 -19+16z=-19 d) y-z=-13 9y+0z=-19 I NEED HELP, ITS REALLY URGENT !!

OpenStudy (amistre64):

whats x when we isolate it?

OpenStudy (anonymous):

wow....

OpenStudy (anonymous):

yeah, basically, just isolate x...

OpenStudy (amistre64):

2(6y-3z)-5+8z=7 12y -6z -5'y' +8z = 7 7y ..... id pick b

OpenStudy (anonymous):

r u sure :( ?

OpenStudy (anonymous):

amistre64 is the Guru, I'm pretty sure his answer is right

OpenStudy (amistre64):

i had to fill in the missing pieces, but yeah :)

OpenStudy (anonymous):

So you're more comfortable with it now niti?

OpenStudy (anonymous):

By subtracting equation [1] from equation [2] you will get b) 7y+2z=7 13y-7z=13